OpenStudy (anonymous):

HURRY!!!!!!! Find a model for simple harmonic motion if the position at t = 0 is 0, the amplitude is 5 centimeters, and the period is 4 seconds.

OpenStudy (dumbcow):

im sorry , is this a race :)

OpenStudy (anonymous):

@dumbcow , the paper is going to finish I suppose :P

OpenStudy (dumbcow):

better guess then...if its multiple choice always choose C

OpenStudy (anonymous):

\[5\sin(\frac{\pi}{2}t)\] is my guess. but i could be wrong

OpenStudy (anonymous):

@satellite73 , you are right :)

OpenStudy (dumbcow):

general form y = Asin(Bt) A is amplitude period = 2pi/B

OpenStudy (anonymous):

do i win?

OpenStudy (anonymous):

Thank you!